# StructEval: Benchmarking LLMs’ Capabilities to Generate Structural Outputs

### v1

@@ Abstract
As Large Language Models (LLMs) become integral to software development workflows, their ability to generate structured outputs has become critically important. We introduce StructEval, a comprehensive benchmark for evaluating LLMs’ capabilities in producing both non-renderable (JSON, YAML, CSV) and renderable (HTML, React, SVG) structured formats. Unlike prior benchmarks, StructEval systematically evaluates structural fidelity across diverse formats through two paradigms: (1) generation tasks, producing structured output from natural language prompts, and (2) conversion tasks, translating between structured formats. Our benchmark encompasses 18 formats and 44 types of task, with novel metrics for format adherence and structural correctness. Results reveal significant performance gaps—even state-of-the-art models like o1-mini achieve only $75.58$ average score, with open-source alternatives lagging approximately $10$ points behind. We find generation tasks more challenging than conversion tasks, and producing correct visual content more difficult than generating text-only structures.
StructEval: Benchmarking LLMs’ Capabilities to Generate Structural Outputs
Jialin Yang $^{*\,\dagger\,1}$  Dongfu Jiang $^{*\,\dagger\,1}$  Lipeng He1  Sherman Siu1  Yuxuan Zhang2 Disen Liao1  Zhuofeng Li4  Huaye Zeng1  Yiming Jia1  Haozhe Wang3 Benjamin Schneider1  Chi Ruan5  Wentao Ma1  Zhiheng Lyu1  Yifei Wang1 Yi Lu2  Quy Duc Do1  Ziyan Jiang1  Ping Nie5  Wenhu Chen† 6 1University of Waterloo, 2University of Toronto, 3HKUST, 4Shanghai University, 5Independent Contributor, 6Vector Institute Equal Contribution {j586yang, dongfu.jiang, wenhuchen}@uwaterloo.ca
@@ 1 Introduction
In recent years, there has been a significant surge in the capabilities of large language models (LLMs) in generating human-like text and performing a wide range of natural language processing tasks. State-of-the-art models like GPT-4o (Hurst et al., 2024), OpenAI o1/o3 (Contributors et al., 2024), and Google’s Gemini (Team et al., 2023) have achieved superior performance in knowledge QA (Hendrycks et al., 2020; Wang et al., 2024), instruction-following (Chiang et al., 2024; Zhou et al., 2023), and code generation (Zhuo et al., 2024; Jain et al., 2024).
Despite recent advances, many real-world applications require not only fluency in the content of the output but also precise control over its structure. This includes tasks where the expected output must follow specific formats such as JSON, XML, LaTeX, HTML, or code in frameworks like React or Vue. Additionally, in these tasks, in these tasks, we also want the code to render a page that correctly places elements according to the requirements. These types of structured output are essential in domains like software development, data pipelines, user interface generation, and scientific publishing, where incorrect formatting can lead to disrupted pipelines or non-functional outputs.
@@
However, most existing benchmarks focus on the semantic quality (Wang et al., 2024) or reasoning ability of LLMs (Hendrycks et al., 2021; He et al., 2024), with limited emphasis on their ability to produce format-conforming structured outputs. Some recently proposed benchmarks aim to evaluate the quality of structured outputs tend to target specific modalities, such as code generation (Zhuo et al., 2024) or text-only structures (Gu et al., 2024; Tang et al., 2023), rather than offering comprehensive evaluations across diverse structured formats. As existing benchmarks gradually become more saturated, it is still unknown how the current state-of-the-art models perform in structured generation tasks. We argue that effectively evaluating the models’ performance on such tasks is inherently challenging due to the following issues:
@@
To address these challenges, we introduce StructEval, a comprehensive benchmark that systematically evaluates LLMs’ abilities to produce highly structured output. Our benchmark encompasses 21 distinct formats and 44 task types organized into two complementary subsets: StructEval-T, which assesses the generation of text-only structures such as JSON and TOML, and StructEval-V, which evaluates the quality of visually rendered outputs from code such as HTML and SVG. Both subsets include generation tasks (converting natural language to structured outputs) and conversion tasks (transforming between two structured formats). To ensure robust evaluation across these diverse formats, we have developed a novel assessment framework that integrates syntactic validity checking, keyword matching, and visual question answering, providing a holistic measure of both structural correctness and output fidelity.
@@ 2.1 Overview
StructEval-T Question, KeyWords Please output JSON code. Task:
@@
StructEval-V Question, Keywords Matching, VQA Pairs Please output HTML code. Task:
@@
Formally, each example is represented as a triplet $(q,\mathbf{K},\mathbf{Q^{v}})$, where $q$ denotes the structure generation question, $\mathbf{K}=\{k_{1},\dots,k_{|\mathbf{K}|}\}$ is a set of keywords expected to appear in the output, and $\mathbf{Q^{v}}=\{(q^{v}_{1},a^{v}_{1}),\dots,(q^{v}_{|\mathbf{Q^{v}}|},a^{v}_{|\mathbf{Q^{v}}|})\}$ is a set of visual question-answer (VQA) pairs used for evaluating examples in the StructEval-V subset. In contrast, for StructEval-T, $\mathbf{Q^{v}}$ is empty and not used during evaluation. To ensure comprehensive evaluation, each example in the dataset contains on average 14.7 keywords and 8.5 VQA pairs, as detailed in Table 1.
@@ 2.2 Annotation Pipeline
To construct a high-quality and diverse benchmark, we design a multi-stage annotation pipeline consisting of three key components: 1) task curation, 2) LLM-based synthesis, and 3) expert review. This pipeline ensures both the scalability and accuracy of the StructEval dataset.
@@
To ensure quality and correctness, we employ a two-pass human review process. Annotators first validate and refine the generated task queries and associated metrics. They are allowed to freely modify, add, or remove any part of the synthesized content to ensure task clarity, completeness, and evaluability. In the second pass, a separate reviewer verifies the consistency and correctness of each example. All annotation is conducted using LabelStudio (Tkachenko et al., 2020-2025), an open-source collaborative annotation tool designed for structured data. The final dataset contains 2035 curated examples, carefully reviewed to support robust evaluation across both StructEval-T and StructEval-V settings.
@@ 3 StructEval Evaluation
VQA Prompt Template You are given an image and a list of question-answer pairs. For each pair, verify if the image content supports the expected answer based on the corresponding question.
@@
This score is used exclusively for tasks in the StructEval-V subset, where the output is expected to be visually rendered. After rendering the output, GPT-4.1-mini (Hurst et al., 2024), a vision-language model (VLM), is employed to answer a set of visual questions $\mathbf{Q^{v}}=\{(q^{v}_{1},a^{v}_{1}),\dots,(q^{v}_{|\mathbf{Q^{v}}|},a^{v}_{|\mathbf{Q^{v}}|})\}$. The VLM will be given both the questions and answers and required to decide whether the VQA pair matches this rendered image. The VQA score is computed as the proportion of correctly answered questions.
Final task scores are calculated as weighted combinations of these metrics, with weights adjusted based on whether the task is renderable. Let $s_{s},s_{k},s_{v}\in[0,1]$ denotes the syntax, keyword matching, and VQA score respectively. The for StructEval-T task, the final score $s$ is computed as:
@@
For StructEval-V, the final score $s$ in computed as:
@@ 4.1 Experimental Setup
We evaluate a range of open-source and commercial large language models (LLMs) using our benchmark. For open-source models, we use Meta-Llama-3-8B-Instruct Grattafiori et al. (2024), Phi-3-mini-128k-instruct Abdin et al. (2024a), Phi-4-mini-instruct Abdin et al. (2024b), Qwen2.5-7B-Instruct Yang et al. (2024), and Qwen3-4B Yang et al. (2025). For commercial models, we use Gemini-1.5-pro and Gemini-2.0-flash Team et al. (2023), GPT-4.1-mini and GPT-4o Hurst et al. (2024), GPT-4o-mini, and o1-mini Contributors et al. (2024). All tasks are evaluated in a zero-shot setting using consistent prompts and parameters.
@@
All model generations are performed using LLM-Engine Jiang (2024), a unified inference framework that supports both open-source backends (e.g., VLLM, SGLang, Together), and commercial APIs (e.g., OpenAI, Claude, Gemini). For open-source models, we specifically utilize the vLLM engine for efficiency  Kwon et al. (2023). For close-source models, we simply call the APIs. As shown in  Table 4, we use greedy decoding by default. All tasks are evaluated zero-shot using uniform task prompts defined in  Table 3. When performing the VQA evaluation, we select GPT-4.1-mini as the VLM due to its superior multimodal abilities (OpenAI, 2025). We apply the VQA prompt template defined in  Figure 5 and ask the VLM to decide whether each VQA pair matches the rendered visual image at once.
@@
Evaluation. Output generations are automatically scored using the evaluation pipeline described in  section 3, including syntactic validity checking, keyword matching, and VQA accuracy. GPT-4.1-mini  (Hurst et al., 2024) is used as the vision-language model for all VQA-based evaluations.
@@ 4.2 Main Results
Llama-3.1-8B-Instruct Grattafiori et al. (2024) 60.22 71.26 54.44 61.15 61.77
Meta-Llama-3-8B-Instruct Grattafiori et al. (2024) 49.18 53.65 46.61 56.91 51.59
Phi-3-mini-128k-instruct Abdin et al. (2024a) 47.39 29.78 44.77 41.23 40.79
Phi-4-mini-instruct Abdin et al. (2024b) 51.38 72.39 51.62 52.48 56.97
Qwen2.5-7B-Instruct  Team (2024) 59.21 62.18 53.28 61.43 59.03
Qwen3-4B  Yang et al. (2025) 64.95 81.13 57.00 65.08 67.04
@@
Gemini-1.5-pro Team et al. (2023) 88.07 74.24 58.11 66.59 71.75
Gemini-2.0-flash Team et al. (2023) 72.42 72.20 53.62 51.97 62.55
GPT-4.1-mini OpenAI (2025) 92.57 75.63 64.30 70.04 75.64
GPT-4o Hurst et al. (2024) 91.52 73.95 65.39 73.20 76.02
GPT-4o-mini Hurst et al. (2024) 79.86 75.57 60.77 76.54 73.19
o1-mini Contributors et al. (2024) 88.12 81.82 61.98 70.40 75.58
@@
Figure 6: Average score over all models based on the most challenging subtasks
@@
As shown in Figure 7, a comparison between generation and conversion tasks in both StructEval-T and StructEval-V settings reveals that, in general, models perform better on conversion tasks than on generation tasks. An exception to this trend occurs in the StructEval-T setting, where commercial models tend to outperform on generation tasks, while open-source models show the opposite behavior—achieving higher scores on conversion tasks.
@@
Figure 7: Average score over all models based on the four task types
@@ 5.1 Large Language Models
Large Language Models (LLMs) have demonstrated remarkable capabilities and gained surging popularity in recent years, ever since the release of ChatGPT (OpenAI, 2023). Over the years, open-source models like Llama (Grattafiori et al., 2024), Phi (Abdin et al., 2024b, a), and Qwen (Yang et al., 2024, 2025) developed by companies like Meta, Microsoft, and Alibaba further facilitated a widespread integration of AI into diverse workflows and everyday applications. Leveraging their large parameter sizes and extensive post-training, LLMs are capable of performing a diverse array of Natural Language Processing (NLP) tasks (Wan et al., 2023). One of the key aspects of the generative capabilities of these models is their ability to generate structured data and transform data from one type to another while maintaining strict adherence to specified formats (Guo et al., 2024). In this paper, we design a new and comprehensive benchmark that evaluates the capability of LLMs to understand, generate, and manipulate structured data across a range of complex, real-world tasks.
@@ 5.2 Evaluation of LLMs
Evaluating structured output has become a focal point for understanding LLM’s limitations (Ning et al., 2025). SoEval (Liu et al., 2024) offers a fast, rule-based check for JSON and XML, but its flat schemas fail to reveal errors in deeper hierarchies. StrucText-Eval (Gu et al., 2024) shifts the task to reasoning over structure-rich text (JSON, YAML, LaTeX) rather than generating the structures themselves, while FOFO (Xia et al., 2024) extends to domains such as law and finance yet covers only a few formats and still relies on human verification. Developer-focused suites like StackEval (Shah et al., 2024) for HTML, CSS, and plotting libraries, and CodeXGLUE (Lu et al., 2021) for multilingual code tasks remain limited to programming artifacts, and Struc-Bench (Tang et al., 2023) concentrates on tabular generation with bespoke metrics. Each benchmark highlights a part of the challenge—be it format adherence, domain coverage, or table fidelity. However, none simultaneously demands broad format coverage, automated grading, and robust transformation capabilities. StructEval addresses these gaps by spanning 18 code and non-code formats, unifying generation, completion, and conversion tasks, and scoring outputs with fully automated structural and vision-based metrics, offering a comprehensive lens on how well LLMs respect and manipulate complex schemas.

### v2

@@ Abstract
As Large Language Models (LLMs) become integral to software development workflows, their ability to generate structured outputs has become critically important. We introduce StructEval, a comprehensive benchmark for evaluating LLMs’ capabilities in producing both non-renderable (JSON, YAML, CSV) and renderable (HTML, React, SVG) structured formats. Unlike prior benchmarks, StructEval systematically evaluates structural fidelity across diverse formats through two paradigms: 1) generation tasks, producing structured output from natural language prompts, and 2) conversion tasks, translating between structured formats. Our benchmark encompasses 18 formats and 44 types of task, with novel metrics for format adherence and structural correctness. Results reveal significant performance gaps—even state-of-the-art models like o1-mini achieve only $75.58$ average score, with open-source alternatives lagging approximately $10$ points behind. We find generation tasks more challenging than conversion tasks, and producing correct visual content more difficult than generating text-only structures. Please see our project page for more detail https://tiger-ai-lab.github.io/StructEval/.
@@ 1 Introduction
In recent years, there has been a significant surge in the capabilities of large language models (LLMs) in generating human-like text and performing a wide range of natural language processing tasks. State-of-the-art models like GPT-4o (Hurst2024GPT4oSC), OpenAI o1/o3 (Contributors2024OpenAIOS), and Google’s Gemini (team2023gemini) have achieved superior performance in knowledge QA (Hendrycks2020MeasuringMM; Wang2024MMLUProAM), instruction-following (Chiang2024ChatbotAA; Zhou2023InstructionFollowingEF), and code generation (Zhuo2024BigCodeBenchBC; Jain2024LiveCodeBenchHA).
Despite recent advances, many real-world applications require not only fluency in the content of the output but also precise control over its structure. This includes tasks where the expected output must follow specific formats such as JSON, XML, LaTeX, HTML, or code in frameworks like React or Vue. Additionally, in these tasks, we also want the code to render a page that correctly places elements according to the requirements. These types of structured output are essential in domains like software development, data pipelines, user interface generation, and scientific publishing, where incorrect formatting can lead to disrupted pipelines or non-functional outputs.
@@
However, most existing benchmarks focus on the semantic quality (Wang2024MMLUProAM) or reasoning ability of LLMs (Hendrycks2021MeasuringMP; He2024OlympiadBenchAC), with limited emphasis on their ability to produce format-conforming structured outputs. Some recently proposed benchmarks aim to evaluate the quality of structured outputs tend to target specific modalities, such as code generation (Zhuo2024BigCodeBenchBC) or text-only structures (gu2024structext; tang-etal-2024-struc), rather than offering comprehensive evaluations across diverse structured formats. As existing benchmarks gradually become more saturated, it is still unknown how the current state-of-the-art models perform in structured generation tasks. We argue that effectively evaluating the models’ performance on such tasks is inherently challenging due to the following issues:
@@
To address these challenges, we introduce StructEval, a comprehensive benchmark that systematically evaluates LLMs’ abilities to produce highly structured output. Our benchmark encompasses 21 distinct formats and 44 task types organized into two complementary subsets: StructEval-T, which assesses the generation of text-only structures such as JSON and TOML, and StructEval-V, which evaluates the quality of visually rendered outputs from code such as HTML and SVG. Both subsets include generation tasks (converting natural language to structured outputs) and conversion tasks (transforming between two structured formats), See Figure 1 for example formats. To ensure robust evaluation across these diverse formats, we have developed a novel assessment framework that integrates syntactic validity checking, keyword matching, and visual question answering, providing a holistic measure of both structural correctness and output fidelity.
@@ 2.1 Overview
StructEval-T Question, KeyWords Please output JSON code.
Task:
@@
StructEval-V Question, Keywords Matching, VQA Pairs Please output HTML code.
Task:
@@
Human Evaluation of VQA Questions Unfair Fair Total Fair Proportion (%)
Correct 6 347 352 98.58%
Wrong 39 6 45 13.33%
Total 44 353 397 88.92%
Accuracy (%) 13.64% 98.30% 88.66%
Table 3: Human evaluation results of sampled VQA questions used in StructEval-V. Each question is annotated as fair or unfair, and correctness is measured by VLM judge performance.
Formally, each example is represented as a triplet $(q,\mathbf{K},\mathbf{Q^{v}})$, where $q$ denotes the structure generation question, $\mathbf{K}=\{k_{1},\dots,k_{|\mathbf{K}|}\}$ is a set of keywords expected to appear in the output, and $\mathbf{Q^{v}}=\{(q^{v}_{1},a^{v}_{1}),\dots,(q^{v}_{|\mathbf{Q^{v}}|},a^{v}_{|\mathbf{Q^{v}}|})\}$ is a set of visual question-answer (VQA) pairs used for evaluating examples in the StructEval-V subset(An example StructEval-V task with keywords and VQA pairs is shown in Figure 4). In contrast, for StructEval-T, $\mathbf{Q^{v}}$ is empty and not used during evaluation (An example StructEval-T question and its keywords are shown in Figure 3). To ensure comprehensive evaluation, each example in the dataset contains on average 14.7 keywords and 8.5 VQA pairs, as detailed in Table 1.
To further assess the quality and fairness of the VQA pairs used in StructEval-V, we conduct a human expert evaluation. Each VQA question is judged as either fair, meaning it can be reasonably answered by a VLM judge using only the rendered image, or unfair, typically involving information not visually accessible, such as precise numeric values or interactive UI elements. Table 3 presents the results of this evaluation. Among 397 sampled VQA pairs, 88.92% were considered fair, and 98.58% of the correct VQA questions were judged fair. Overall, 98.30% of all fair questions could be correctly answered by our VLM judge (GPT-4.1-mini), supporting the validity of our automated evaluation process.
@@ 2.2 Annotation Pipeline
To construct a high-quality and diverse benchmark, we design a multi-stage annotation pipeline consisting of three key components: 1) task curation, 2) LLM-based synthesis, and 3) expert review (see Figure 2 for an overview of this pipeline). This pipeline ensures both the scalability and accuracy of the StructEval dataset.
@@
To ensure quality and correctness, we employ a two-pass human review process. Annotators first validate and refine the generated task queries and associated metrics. They are allowed to freely modify, add, or remove any part of the synthesized content to ensure task clarity, completeness, and evaluability. In the second pass, a separate reviewer verifies the consistency and correctness of each example. All annotation is conducted using LabelStudio (labelstudio), an open-source collaborative annotation tool designed for structured data. The final dataset contains 2035 curated examples, carefully reviewed to support robust evaluation across both StructEval-T and StructEval-V settings.
@@ 3 StructEval Evaluation
VQA Prompt Template You are given an image and a list of question-answer pairs.
For each pair, verify if the image content supports the expected answer based on the corresponding question.
@@
This score is used exclusively for tasks in the StructEval-V subset, where the output is expected to be visually rendered. After rendering the output, GPT-4.1-mini (Hurst2024GPT4oSC), a vision-language model (VLM), is employed to answer a set of visual questions $\mathbf{Q^{v}}=\{(q^{v}_{1},a^{v}_{1}),\dots,(q^{v}_{|\mathbf{Q^{v}}|},a^{v}_{|\mathbf{Q^{v}}|})\}$. The VLM will be given both the questions and answers and required to decide whether the VQA pair matches this rendered image. The VQA score is computed as the proportion of correctly answered questions.
Final task scores are calculated as weighted combinations of these metrics, with weights adjusted based on whether the task is renderable. Let $s_{s},s_{k},s_{v}\in[0,1]$ denotes the syntax, keyword matching, and VQA score respectively. For the StructEval-T task, the final score $s$ is computed as:
@@
For StructEval-V, the final score $s$ is computed as:
@@ 4.1 Experimental Setup
We evaluate a range of open-source and commercial large language models (LLMs) using our benchmark. For open-source models, we use Meta-Llama-3-8B-Instruct grattafiori2024llama3herdmodels, Phi-3-mini-128k-instruct abdin2024phi3technicalreporthighly, Phi-4-mini-instruct abdin2024phi4technicalreport, Qwen2.5-7B-Instruct yang2024qwen2technicalreport, and Qwen3-4B yang2025qwen3technicalreport. For commercial models, we use Gemini-1.5-pro and Gemini-2.0-flash team2023gemini, GPT-4.1-mini and GPT-4o Hurst2024GPT4oSC, GPT-4o-mini, and o1-mini Contributors2024OpenAIOS. All tasks are evaluated in a zero-shot setting using consistent prompts and parameters.
@@
All model generations are performed using LLM-Engine jiang2024llmengines, a unified inference framework that supports both open-source backends (e.g., VLLM, SGLang, Together), and commercial APIs (e.g., OpenAI, Claude, Gemini). For open-source models, we specifically utilize the vLLM engine for efficiency  kwon2023efficient. For closed source models, we simply call the APIs. As shown in  Table 5, we use greedy decoding by default. All tasks are evaluated zero-shot using uniform task prompts defined in  Table 4. When performing the VQA evaluation, we select GPT-4.1-mini as the VLM due to its superior multimodal abilities (openai2025gpt41). We apply the VQA prompt template defined in  Figure 5 and ask the VLM to decide whether each VQA pair matches the rendered visual image at once.
@@
Evaluation. Output generations are automatically scored using the evaluation pipeline described in  section 3, including syntactic validity checking, keyword matching, and VQA accuracy. GPT-4.1-mini  (Hurst2024GPT4oSC) is used as the vision-language model for all VQA-based evaluations.
@@ 4.2 Main Results
Llama-3.1-8B-Instruct grattafiori2024llama3herdmodels 60.22 71.26 54.44 61.15 61.77
Meta-Llama-3-8B-Instruct grattafiori2024llama3herdmodels 49.18 53.65 46.61 56.91 51.59
Phi-3-mini-128k-instruct abdin2024phi3technicalreporthighly 47.39 29.78 44.77 41.23 40.79
Phi-4-mini-instruct abdin2024phi4technicalreport 51.38 72.39 51.62 52.48 56.97
Qwen2.5-7B-Instruct  qwen2.5 59.21 62.18 53.28 61.43 59.03
Qwen3-4B  yang2025qwen3technicalreport 64.95 81.13 57.00 65.08 67.04
@@
Gemini-1.5-pro team2023gemini 88.07 74.24 58.11 66.59 71.75
Gemini-2.0-flash team2023gemini 72.42 72.20 53.62 51.97 62.55
GPT-4.1-mini openai2025gpt41 92.57 75.63 64.30 70.04 75.64
GPT-4o Hurst2024GPT4oSC 91.52 73.95 65.39 73.20 76.02
GPT-4o-mini Hurst2024GPT4oSC 79.86 75.57 60.77 76.54 73.19
o1-mini Contributors2024OpenAIOS 88.12 81.82 61.98 70.40 75.58
@@
(a) Avg. score over all models based on most challenging subtasks
(b) Avg. score over all models based on the four task types
@@
As shown in 6(b), a comparison between generation and conversion tasks in both StructEval-T and StructEval-V settings reveals that, in general, models perform better on conversion tasks than on generation tasks. An exception to this trend occurs in the StructEval-T setting, where commercial models tend to outperform on generation tasks, while open-source models show the opposite behavior—achieving higher scores on conversion tasks.
@@ 5.1 Large Language Models
Large Language Models (LLMs) have demonstrated remarkable capabilities and gained surging popularity in recent years, ever since the release of ChatGPT (OpenAI2023ChatGPT). Over the years, open-source models like Llama (grattafiori2024llama3herdmodels), Phi (abdin2024phi4technicalreport; abdin2024phi3technicalreporthighly), and Qwen (yang2024qwen2technicalreport; yang2025qwen3technicalreport) developed by companies like Meta, Microsoft, and Alibaba further facilitated a widespread integration of AI into diverse workflows and everyday applications. Leveraging their large parameter sizes and extensive post-training, LLMs are capable of performing a diverse array of Natural Language Processing (NLP) tasks (wan2023gptreincontextlearningrelation). One of the key aspects of the generative capabilities of these models is their ability to generate structured data and transform data from one type to another while maintaining strict adherence to specified formats (guo2024large). In this paper, we design a new and comprehensive benchmark that evaluates the capability of LLMs to understand, generate, and manipulate structured data across a range of complex, real-world tasks.
@@ 5.2 Evaluation of LLMs
Evaluating structured output has become a focal point for understanding LLM’s limitations (ning2025picopeerreviewllms). SoEval (liu2024soeval) offers a fast, rule-based check for JSON and XML, but its flat schemas fail to reveal errors in deeper hierarchies. StrucText-Eval (gu2024structext) shifts the task to reasoning over structure-rich text (JSON, YAML, LaTeX) rather than generating the structures themselves, while FOFO (xia2024fofo) extends to domains such as law and finance yet covers only a few formats and still relies on human verification. Developer-focused suites like StackEval (shah2024stackeval) for HTML, CSS, and plotting libraries, and CodeXGLUE (lu2021codexglue) for multilingual code tasks remain limited to programming artifacts, and Struc-Bench (tang-etal-2024-struc) concentrates on tabular generation with bespoke metrics. Each benchmark highlights a part of the challenge—be it format adherence, domain coverage, or table fidelity. However, none simultaneously demands broad format coverage, automated grading, and robust transformation capabilities. StructEval addresses these gaps by spanning 18 code and non-code formats, unifying generation, completion, and conversion tasks, and scoring outputs with fully automated structural and vision-based metrics, offering a comprehensive lens on how well LLMs respect and manipulate complex schemas.